# MHNet: Multi-view High-order Network for Diagnosing Neurodevelopmental Disorders Using Resting-state fMRI

### v1

@@ 5.2.2 Influence of Brain Atlas Selection
Figure 4: Comparision of the classification performance between using AAL1 and Brainnetome.
@@ 5.2.4 Influence of Multi-view and High-order Feature Extractors
Table 4: Classification performance for ASD (ASD vs. NC) classification results on ABIDE I and ABIDE II, and ADHD (ADHD vs. NC) classification results on ADHD-200 using varied multi-view data and advanced features. Percentages represent the outcomes from 10-fold cross-validation. The best result in each category is highlighted in bold.
@@
ABIDE-I GNN(Only Brain-LAN)1 63.39 ± 2.59 65.15 ± 1.53
GNN2 66.62 ± 0.76 67.37 ± 2.54
HGNN3 71.48 ± 1.23 73.35 ± 1.46
HGNN + CNN4 73.06 ± 2.03 74.58 ± 1.03
HGNN + HCNN5 76.29 ± 1.39 78.68 ± 2.07
ABIDE-II GNN(Only Brain-LAN) 64.45 ± 1.03 66.28 ± 0.93
GNN 65.31 ± 2.03 67.15 ± 1.74
HGNN 69.49 ± 1.71 71.23 ± 0.99
HGNN + CNN 71.35 ± 1.43 73.53 ± 2.23
HGNN + HCNN 75.16 ± 1.63 76.28 ± 1.82
ADHD-200 GNN(Only Brain-LAN) 62.18 ± 2.96 65.29 ± 1.37
GNN 62.03 ± 2.23 63.48 ± 1.27
HGNN 67.12 ± 0.74 68.23 ± 1.25
HGNN + CNN 68.54 ± 2.37 70.34 ± 1.56
HGNN + HCNN 70.33 ± 0.76 72.02 ± 0.86
@@
1 GNN (Only Brain-LAN) means that only the Brian-LAN layer in G-IBHN-G is used to obtain brain FC features and high-order features and HCNN branches are not used; 2 GNN means that a complete BFN with hierarchical structure in G-IBHN-G is used, but high-order features and HCNN branches are not used; 3HGNN uses high-order feature representation on the basis of GNN; 4 HGNN + CNN means that the FC information of the brain in Euclidean space is also used on the basis of HGNN, but high-order features are not used; 5 HGNN + HCNN is the complete MHNet model.
@@
As shown in Table 4, without encoding high-order features and multi-view information, GNN (Only Brain-LAN) demonstrates basic performance, and all evaluation indicators are relatively low. Compared with GNN (Only Brain-LAN), the performance of GNN is improved, which means that the hierarchical structure of BFN can capture more useful features. Based on GNN, HGNN is proposed to encode high-order features, which significantly improves the performance of the model, especially in capturing the complex FC of the brain. Apart from HGNN, we also add CNN which encodes the FC information of the brain in Euclidean space, resulting in a HGNN + CNN model. Although this model introduces the Euclidean information in FC, the performance improvement is not as significant as that of HGNN. The high-order features may play a more important role in capturing complex activity patterns of the brain. HGNN + HCNN shows the best performance with respect to all evaluation indicators, indicating the effectiveness of introducing the complementary multi-view high-order features.
@@ 6.1 Interpretability Analysis
ASD involves abnormalities in multiple brain regions, including the lingual gyrus, emporal inferior gyrus, hippocampus, amygdala, calcarine fissure, and occipital middle gyrus [58]. The lingual gyrus plays an important role in visual processing, especially in facial recognition and emotional understanding. The abnormality of lingual gyrus may lead to difficulties in visual information integration and social interaction for patients with ASD [59]. The inferior temporal gyrus and the temporal middle gyrus play an important role in language processing and social cognition, and their functional abnormalities may explain the language comprehension and episodic memory impairments for patients with ASD [60]. Abnormalities in the hippocampus and parahippocampal gyrus may lead to memory and spatial cognition problems. While abnormalities in amygdala are associated with the problems in emotional processing and social behavior regulation [61]. Abnormalities in the calcarine fissure and the occipital middle gyrus may affect primary and higher-order processing of visual information, which explains the challenges that ASD patients face in processing visual social cues [62].
In addition, regions such as the insula, precuneus, fusiform gyrus, parietal superior lobule, and caudate nucleus also show significant abnormalities in ASD patients. The insula plays an important role in emotion processing and interoception, and its functional changes may lead to difficulties in emotion understanding and self-awareness for ASD patients [63]. The precuneus is related to self-reflection and social cognition, and its abnormalities may affect self-awareness and understanding of others in social interactions. The fusiform gyrus plays an important role in facial recognition, and its abnormalities may lead to difficulties in ASD patients in recognizing faces and interpreting facial expressions. Abnormalities in the superior parietal lobule may affect spatial cognition and visual-spatial processing, which explains the challenges of ASD patients in spatial navigation and task switching. The caudate nucleus plays an important role in motor control, learning, and executive function, and its functional abnormalities may be related to repetitive behaviors, narrow interests, and executive dysfunction in ASD patients [64]. The combined changes in these brain regions lead to a wide range of challenges in cognitive, emotional, and social functions for ASD patients.
@@ 6.3 Non-imaging Data
In this work, our MHNet model achieves significant diagnostic results. However, our proposed MHNet model does not take into accout the non-imaging data. The phenotypic data, such as clinical symptoms and behavioral assessments, play an important role in NDD diagnosis. Fusing non-imaging data into our current framework is an interesting direction that will be explored to improve its diagnostic performance in future study. By integrating phenotypic information [70], models can capture more pathological features and individual differences, resulting in more accurate diagnosis performance.

### v2

@@ 5.2.2 Influence of Brain Atlas Selection
Figure 5: Comparision of the classification performance between using AAL1 and Brainnetome. * indicates a statistically significant differences.
@@ 5.2.4 Influence of Multi-view and High-order Feature Extractors
Table 4: Classification performance for ASD (ASD vs. NC) classification results on ABIDE I and ABIDE II, and ADHD (ADHD vs. NC) classification results on ADHD-200 using varied multi-view data and advanced features. Percentages represent the outcomes from 10-fold cross-validation. Statistically significant differences from MHNet marked *. The best result in each category is highlighted in bold.
@@
ABIDE-I GNN(Only Brain-LAN)1 63.39 ± 2.59* 65.15 ± 1.53*
GNN2 66.62 ± 0.76* 67.37 ± 2.54*
CNN3 64.39 ± 1.85* 65.93 ± 1.36*
HGNN4 71.48 ± 1.23* 73.35 ± 1.46*
HCNN5 70.19 ± 2.04* 72.66 ± 1.37*
HCNN + GNN6 71.93 ± 2.29* 73.02 ± 4.32
HGNN + CNN7 73.06 ± 2.03 74.58 ± 1.03*
HGNN + HCNN8 76.29 ± 1.39 78.68 ± 2.07
ABIDE-II GNN(Only Brain-LAN)1 64.45 ± 1.03* 66.15 ± 1.53*
GNN2 65.31 ± 2.03* 67.15 ± 1.74*
CNN3 65.91 ± 1.54* 65.19 ± 1.49*
HGNN4 69.49 ± 1.71* 71.23 ± 0.99*
HCNN5 68.45 ± 3.13* 70.22 ± 1.95*
HCNN + GNN6 70.59 ± 1.62* 73.28 ± 1.57*
HGNN + CNN7 71.35 ± 1.43* 73.53 ± 2.23
HGNN + HCNN8 75.16 ± 1.63 76.28 ± 1.82
ADHD-200 GNN(Only Brain-LAN)1 62.18 ± 2.96* 65.29 ± 1.37*
GNN2 62.03 ± 2.23* 63.48 ± 1.27*
CNN3 63.26 ± 1.59* 64.03 ± 1.93*
HGNN4 67.12 ± 0.74* 68.23 ± 1.25*
HCNN5 67.38 ± 2.19* 67.69 ± 1.73*
HCNN + GNN6 68.05 ± 1.21* 69.59 ± 2.24*
HGNN + CNN7 68.54 ± 2.37 70.34 ± 1.56*
HGNN + HCNN8 70.33 ± 0.76 72.02 ± 0.86
@@
11footnotetext: GNN (Only Brain-LAN) means that only the Brian-LAN layer in G-IBHN-G is used to obtain brain FC features and high-order features and HCNN branches are not used.22footnotetext: GNN means that a complete BFN with hierarchical structure in G-IBHN-G is used, but high-order features and HCNN branches are not used.
33footnotetext: CNN means that a complete BFN with hierarchical structure in G-IBHN-G is used, but high-order features and HCNN branches are not used.44footnotetext: HGNN uses high-order feature representation on the basis of GNN.
55footnotetext: HCNN uses high-order feature representation on the basis of CNN.66footnotetext: HCNN + GNN means that the FC information of the brain in Euclidean space is also used on the basis of HGNN, but high-order features are not used.
77footnotetext: HGNN + CNN means that the FC information of the brain in Euclidean space is also used on the basis of HGNN, but high-order features are not used.88footnotetext: HGNN + HCNN is the complete MHNet model.
@@
As shown in Table 4, without encoding high-order features and multi-view information, GNN (Only Brain-LAN) demonstrates basic performance, and all evaluation indicators are relatively low. Compared with GNN (Only Brain-LAN), the performance of GNN is improved, which means that the hierarchical structure of BFN can capture more useful features. Similarly, CNN achieves comparable results, suggesting that Euclidean-based FC features also contribute to the classification performance. Based on GNN, HGNN is proposed to encode high-order features, which significantly improves the performance of the model, especially in capturing the complex FC of the brain. HCNN, which introduces high-order features on the basis of CNN, also shows notable performance enhancement. Combining HCNN with GNN yields better results than using GNN alone, highlighting the benefit of high-order features. HCNN + GNN combines Euclidean high-order features with non-Euclidean network structure, showing moderate improvement compared to single-view methods. Apart from HGNN, we also add CNN which encodes the FC information of the brain in Euclidean space, resulting in a HGNN + CNN model. Although this model introduces the Euclidean information in FC, the performance improvement is not as significant as that of HGNN. The high-order features may play a more important role in capturing complex activity patterns of the brain. HGNN + HCNN shows the best performance with respect to all evaluation indicators, indicating the effectiveness of introducing the complementary multi-view high-order features.
5.2.5 Influence of Network Architecture Parameters
Table 5: Different Network Architecture Parameters on Classification Performance (%). Statistically significant differences from MHNet with optimal parameters marked *. The best results are shown in bold.
Parameter Value ABIDE-I ABIDE-II ADHD-200
ACC AUC ACC AUC ACC AUC
ChebNet Order (K) 1 64.45* 65.89* 63.23* 64.98* 59.12* 61.89*
2 72.78* 73.12* 70.45* 72.87* 67.91* 69.45*
3 76.29 78.68 75.16 76.28 70.33 72.02
4 73.89* 74.12* 72.56* 73.91* 69.17* 70.12*
5 71.12* 71.67* 69.89* 70.45* 65.12* 68.78*
#ChebConv Modules 1 68.12* 69.11* 67.37* 68.03* 63.43* 64.72*
2 73.15* 73.98* 72.32* 73.87* 67.23* 69.29*
3 76.29 78.68 75.16 76.28 70.33 72.02
4 74.38* 75.15* 73.52* 73.73* 68.28* 70.25*
To investigate the impact of key architectural parameters on model performance, we conducted extensive ablation studies focusing on the polynomial order (K) of ChebNet and the number of ChebConv modules in Brainnetome Atlas. For the polynomial order, we varied K from 1 to 5 while keeping other parameters fixed. As shown in Table 5, the model achieves optimal performance with K=3 on all three datasets, with accuracies of 76.29%, 75.16%, and 70.33% for ABIDE-I, ABIDE-II, and ADHD-200 respectively. When K $>$ 3, the performance begins to degrade due to increased model complexity and potential overfitting. Similarly, we examined the effect of ChebConv module depth by varying the number of layers from 1 to 4. The results in Table 5 demonstrate that three ChebConv modules provide the best balance between model capacity and computational efficiency. Adding more layers leads to marginal improvements while significantly increasing computational cost. These findings guided our final architecture design, which uses K=3 and three ChebConv modules.
@@ 6.1 Interpretability Analysis
ASD involves abnormalities in multiple brain regions, including the lingual gyrus, emporal inferior gyrus, hippocampus, amygdala, calcarine fissure, and occipital middle gyrus [63]. The lingual gyrus plays an important role in visual processing, especially in facial recognition and emotional understanding. The abnormality of lingual gyrus may lead to difficulties in visual information integration and social interaction for patients with ASD [64]. The inferior temporal gyrus and the temporal middle gyrus play an important role in language processing and social cognition, and their functional abnormalities may explain the language comprehension and episodic memory impairments for patients with ASD [65]. While abnormalities in amygdala are associated with the problems in emotional processing and social behavior regulation [66]. Abnormalities in the calcarine fissure and the occipital middle gyrus may affect primary and higher-order processing of visual information, which explains the challenges that ASD patients face in processing visual social cues [67]. In addition, the insula, crucial for emotion processing and interoception, along with precuneus, fusiform gyrus, parietal superior lobule, and caudate nucleus show significant abnormalities in ASD patients [68]. The precuneus governs self-reflection and social cognition, fusiform gyrus controls facial recognition, and superior parietal lobule manages spatial cognition, all showing abnormalities impacting ASD patients’ social and cognitive functions. The caudate nucleus plays an important role in motor control, learning, and executive function, and its functional abnormalities may be related to repetitive behaviors, narrow interests, and executive dysfunction in ASD patients [69]. The combined changes in these brain regions lead to a wide range of challenges in cognitive, emotional, and social functions for ASD patients.
@@ 6.2 Brain Feature Analysis
Figure 8: Architecture of population graph incorporating phenotypic information.
@@ 6.3 Non-imaging Data
In this work, our MHNet model achieves significant diagnostic results. When designing MHNet, we focused on designing models that are good at identifying and analyzing local brain regions and biomarkers associated with diseases. However, our proposed MHNet model does not take into accout the non-imaging data. Non-imaging data such as clinical symptoms and behavioral assessments are likely to play an important role in the diagnosis of NDD [75].
To further explore the impact of non-imaging data, we considered the relationships among subjects and tried to exploit phenotypic information such as gender, age, and site. First, leveraging MHNet’s exceptional performance, we employ it to generate precise embeddings that effectively represent brain networks. Then, we construct a population graph for subject classification, with each node representing an individual subject’s brain and each edge capturing the relationship between pairs of subjects. The specific framework of the population graph is shown in Figure 8.
To achieve a comprehensive representation of all subjects $S=\{{S}_{1},{S}_{2},\ldots,{S}_{m}\}$ using both imaging and non-imaging data, we construct a population graph $G_{\text{p}}=\{{V}^{\prime},{A}^{\prime}\}$. Here, $m$ denotes the number of subjects, so the set $V^{\prime}$ contains $m$ nodes. Each node in $V^{\prime}$ corresponds to an embedding $Y$ of a subject $S$, obtained by reshaping the output of the MHNet into a vector form. The adjacency matrix $A^{\prime}$ for the graph is defined as ${A}^{\prime}={C}*{W}$, where $C$ is a binarized connectivity matrix derived from both imaging and non-imaging data. To construct $C$, we first create a similarity matrix ${M}_{1}$ using the embedding features $Y$:
${M}_{1}=\exp\left(-\frac{[\rho({Y}_{i},{Y}_{j})]^{2}}{2\sigma^{2}}\right),$ (19)
where $\rho$ represents the correlation distance between embeddings, and $\sigma$ is the mean of $[\rho({Y}_{i},{Y}_{j})]^{2}$. Next, we develop a node similarity metric ${M}_{2}$ by incorporating phenotypic information (e.g., gender) to obtain more useful clinical auxiliary information. These two matrices are then combined using the Hadamard product to form ${C}^{\prime}={M}_{1}*{M}_{2}$.
Table 6: Classification performance for ASD (ASD vs. NC) classification results on ABIDE I and ABIDE II, and ADHD (ADHD vs. NC) classification results on ADHD-200 using non-imaging data. Percentages represent the outcomes from 10-fold cross-validation. Statistically significant differences from MHNet marked *. The best result in each category is highlighted in bold.
Dataset Method Non-imaging Data ACC AUC
ABIDE-I GCN (Regional Brain Graph)1 $\times$ 66.62 ± 0.76* 67.37 ± 2.54*
GCN (Population Graph)2 $\checkmark$ 69.13 ± 1.24* 69.31 ± 2.38*
MHNet (Regional Brain Graph)3 $\times$ 76.29 ± 1.39* 78.68 ± 2.07
MHNet (Population Graph)4 $\checkmark$ 77.03 ± 2.47 78.51 ± 1.39
ABIDE-II GCN (Regional Brain Graph)1 $\times$ 65.31 ± 2.03* 67.15 ± 1.74*
GCN (Population Graph)2 $\checkmark$ 66.29 ± 1.53* 68.93 ± 1.89*
MHNet (Regional Brain Graph)3 $\times$ 75.16 ± 1.63 76.28 ± 1.82
MHNet (Population Graph)4 $\checkmark$ 75.73 ± 2.02 77.39 ± 1.31
ADHD-200 GCN (Regional Brain Graph)1 $\times$ 62.03 ± 2.23* 63.48 ± 1.27*
GCN (Population Graph)2 $\checkmark$ 63.19 ± 3.04* 63.29 ± 2.69*
MHNet (Regional Brain Graph)3 $\times$ 70.33 ± 0.76* 72.02 ± 0.86*
MHNet (Population Graph)4 $\checkmark$ 72.12 ± 2.17 72.39 ± 1.53
11footnotetext: GCN (Regional Brain Graph): GCN for regional brain graph classification, without phenotypic data.22footnotetext: GCN (Population Graph): GCN with phenotypic data for population graph node classification.
33footnotetext: MHNet (Regional Brain Graph): GCN for regional brain graph classification, without phenotypic data.44footnotetext: MHNet (Population Graph): MHNet to capture node features and construct population graph with phenotypic data.
From $C^{\prime}$, we obtain the binarized connection graph $C$. The sparse weight matrix $W$ is constructed by calculating each element ${W}_{ij}$ based on the normalized non-imaging data ${\eta}_{i}$ and ${\eta}_{j}$ of two subjects:
${W}_{ij}=\frac{Sim\left(\text{MLP}(\eta_{i}),\text{MLP}(\eta_{j})\right)+1}{2},$ (20)
In this equation, the MLP layers share weights, and $Sim$ denotes the cosine similarity between the outputs of the MLPs applied to ${\eta}_{i}$ and ${\eta}_{j}$.
Next, we use GCN and MLP to classify nodes and get the classification result $y$. The specific operations are as follows:
$y=MLP(GCN(Y,A^{\prime})).$ (21)
As shown in Table 6, incorporating non-imaging data through population graphs yields modest improvements in classification performance. While the MHNet with population graph achieves the best performance across all datasets, the improvements over MHNet with regional brain graph are relatively small. This limited improvement can be attributed to our initial focus in MHNet design, which primarily emphasized the identification and analysis of local brain regions and disease-related biomarkers rather than the integration of non-imaging features. Nevertheless, statistical analysis reveals that MHNet-based methods (both with and without population graphs) significantly outperform GNN-based approaches, demonstrating the robustness of MHNet’s brain network feature extraction capabilities. The results suggest that while non-imaging data provides complementary information, the core strength of MHNet lies in its ability to effectively capture disease-specific patterns from brain imaging data.
6.4 Study Limitations
While MHNet demonstrates promising results in NDD diagnosis, several aspects warrant further investigation. Although validated on three public datasets using AAL1 and Brainnetome Atlas templates, incorporating more diverse datasets and brain atlas templates would enhance generalizability across different clinical settings and demographics. The computational complexity increases with multiple views and hierarchical levels, necessitating optimization for real-time clinical applications, while more effective multi-view integration strategies could improve clinical utility. Beyond basic demographics, expanding phenotypic features to include behavioral assessments, cognitive profiles, and environmental factors, coupled with systematic evaluation of their relative contributions to diagnostic accuracy, could provide more nuanced NDD characterization and guide clinical data collection priorities.
@@ 7 Conclusion
Declarations
[…]
Author contribution Conceptualization: Yueyang Li, Weiming Zeng, NiZhuan Wang. Methodology: Yueyang Li, Weiming Zeng, NiZhuan Wang. Software: Yueyang Li, Wenhao Dong, Lei Wang. Validation: Yueyang Li, Weiming Zeng, NiZhuan Wang, Luhui Cai, Hongjie Yan, Lingbin Bian. Formal analysis:, Yueyang Li, Luhui Cai, Wenhao Dong. Investigation: Yueyang Li, Weiming Zeng, NiZhuan Wang, Hongjie Yan, Lingbin Bian, Wenhao Dong, Lei Wang. Data curation: Yueyang Li, Wenhao Dong, Luhui Cai. Writing—original draft preparation: all authors. Writing—review and editing: all authors. Visualization: Yueyang Li, Weiming Zeng, NiZhuan Wang, Wenhao Dong, Luhui Cai. Supervision, Weiming Zeng, NiZhuan Wang. Project administration: Weiming Zeng, NiZhuan Wang. Funding acquisition: Weiming Zeng. All authors have read and agreed to the published version of the manuscript.